OpenStudy (anonymous):

find indefinite integral of csc^2(t)/cot(t) dx plz help!!!

OpenStudy (anonymous):

Remember: (1) d/dt [cot t] = -csc^2 t (2) sin 2x = 2 sin x cos x By (1) we see that INT {d/dt [cot t]} dt = INT {-csc^2 t} dt so cot t = - INT {csc^2 t} dt so INT {csc^2 t} dt = -cot t. a) INT {6(csc t)^2 - 4e^t} dt 6(-cot t) - 4e^t + C b) INT {5 sin 2x / sin x} dx INT {5 * 2sin x cos x / sin x} dx INT {10 cos x} dx 10 sin x + C

OpenStudy (anonymous):

@angelarios did that clear everything for you?

OpenStudy (anonymous):

im a little confused :/ @Gracy_KSL

OpenStudy (anonymous):

what are you confused about?

OpenStudy (anonymous):

I don't know where you're getting the very first part :o @Gracy_KSL

OpenStudy (anonymous):

what do u mean why? or dt? @Loser66

OpenStudy (anonymous):

oh yea sorry I meant dt :p sorry :( @Loser66

OpenStudy (loser66):

makes sense?

OpenStudy (anonymous):

yes but I don't know where you are getting the cos/sin :o isn't it when you multiply 1/sin^2 by cos/sin isn't it cos/sin^3 ??

OpenStudy (loser66):

your problem is \[\large\int \dfrac{csc^2 t}{cot t}dt\] right?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

numerator \(csc ^2t = \dfrac{1}{sin^2t}\) ok?

OpenStudy (anonymous):

ok

OpenStudy (loser66):

denominator \(cot t= \dfrac{cos t}{sin t}\) right?

OpenStudy (anonymous):

yes and after that you do numerator times reciprocal of denominator right??

OpenStudy (loser66):

yes, that's the way I do, but I had some mistake on the process, you can do the next step

OpenStudy (loser66):